OpenStudy (anonymous):

Find the point on the terminal side of θ =–3pi/4 that has an x coordinate of –1.

hero (hero):

1. let pi = 180: θ = -3(180)/4 θ = -3(45) θ = -135 θ = 225 degrees Let x = -1 Now create right triangle: